OpenStudy (anonymous):

medal & fan (: Choose the equation of the horizontal line that passes through the point (1, -5). y = -5 x = 1 y = 1 x = -5

OpenStudy (shamil98):

horizontally on the rectangular coordinate system is the y so y = -5 

OpenStudy (anonymous):

thank you (: will you help me on another ?

OpenStudy (shamil98):

sure.

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (shamil98):

plug in the values of x and y and m into y = mx + b and solve for b. that way you find your y-intercept

OpenStudy (anonymous):

Y=-2 -3+ -6 is that right ?

OpenStudy (shamil98):

Close. remember y = mx + b (-2,-3) <-- (x,y) , m = -6 -3 = -6(-2) + b

OpenStudy (anonymous):

what is b ?

OpenStudy (shamil98):

b is your y intercept where you start graphing the line

OpenStudy (anonymous):

so, -3= -6(-2)+ b .. am i suppose to solve thatt ?

OpenStudy (shamil98):

Yeah solve for b

OpenStudy (anonymous):

how? lol

OpenStudy (shamil98):

-3 = -6(-2) + b -3 = 12 + b -12 -12 -15 = b

OpenStudy (anonymous):

thank you.. I understand now